OpenStudy (anonymous):

For which equation does the point (20,60) lie on the graph ? Explain i. C=15+2.50n ii. C=50+1.75n iii. C=30+1.50n

OpenStudy (anonymous):

Please help

OpenStudy (amistre64):

i assume its in the form (n,C)

OpenStudy (anonymous):

What ?

OpenStudy (amistre64):

iii is my first guess; but would have to dbl check it

OpenStudy (anonymous):

amistre64 check out the method i gav for finding the quadratic equation given its vertex...please :)

OpenStudy (amistre64):

60 = 30 + 1.5(20) 60 = 30 + 30 60 = 60

OpenStudy (anonymous):

Okay , I'm going to check ...

OpenStudy (anonymous):

60,20 only fits no. iii

OpenStudy (amistre64):

the (n,C) refers to the way they expect you to fill in the question; the independant variable tends to be first; then the dependant

OpenStudy (amistre64):

(if I plug in 'n', do I get 'C') ( 20 , 60 )

OpenStudy (anonymous):

Yes, I know thiss .

OpenStudy (anonymous):

Thanks